Tori invested $8000 in money market funds. Part was invested at 5% simple interest and the rest at 7% simple interest. At the end of 1 year, Tori had earned $496 in interest. How much did she invest in each fund?
Answer by stanbon(26291)

Tori invested $8000 in money market funds. Part was invested at 5% simple interest and the rest at 7% simple interest. At the end of 1 year, Tori had earned $496 in interest. How much did she invest in each fund?
--------------------
Interest Equation:
0.05x + 0.07(8000-x) = 496
---
Multiply thru by 100 to get:
5x + 7*8000 - 7x = 49600
-2x = -6400
x = $3200 (amount invested at 5%)
----
8000-3200 = $4800 (amount invested at 7%)
============================================
Cheers,
Stan H.